Circular No. 8144

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1814–338 C. B. Markwardt, University of Maryland and Goddard Space Flight

Center (GSFC); and J. H. Swank, GSFC, report that RXTE PCA monitoring observations of the Galactic-center region have revealed a new millisecond pulsar, designated XTE J1814–338. PCA cross scans on June 5.1 UT determined the position to be α = 18^{h}13^{m}.7, δ = −33^{o}46′ (equinox 2000.0), with a 99-percent confidence radius of 2′. The observation also contained pointed data, during which coherent pulsations were detected at a barycentered frequency of 314.319 Hz, making this the fifth known millisecond accreting pulsar. There appears to be a significant sinusoidal modulation of the pulse frequency, at a period of 1.9 ± 0.3 hr, with a half-amplitude of 0.014 ± 0.002 Hz, although the short time baseline (3400 s) may contribute significant systematic uncertainties. Interpreted as an orbital Doppler modulation, the orbital period is most similar to the first such pulsar, SAX J1808.4–3658 (IAUC 6876, 6877). The following 2–10keV x-ray fluxes (in mCrab, with uncertainties of ≈ 0.1 mCrab) indicate that the source intensity is still increasing: May 27.5, 0.1; 31.5, 1.8; June 3.5, 7.1; 5.1, 10.5. Observations at other wavelengths are encouraged.

SUPERNOVA 2003ed IN NGC 5303A D. C. Leonard, University of Massachusetts, Amherst; and R. Chor-

nock and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 300–930 nm) of the type-II SN 2003ed (IAUC 8129), obtained on May 25 UT with the Keck-I telescope, show it to bear a striking resemblance to the spectrum of the type-IIb SN 1993J, ∼ 32 days past explosion (Matheson et al. 2000, A.J. 120, 1487). In addition to features typically seen in type-II supernovae at this stage, lines of He I with P-Cyg profiles have emerged, suggesting that the progenitor of SN 2003ed lost a substantial fraction of its hydrogen envelope prior to exploding. In addition, spectropolarimetry (range 570–940 nm) shows an average continuum polarization of ∼ 1 percent, with strong modulations across the Hα and Ca II-infrared-triplet features of up to ∼ 1 percent in the individual Stokes parameters. While some of the continuum polarization may be due to interstellar dust, the changes across the line features suggest substantial polarization due to scattering by aspherical supernova ejecta, as has been inferred for SN 1993J at a similar epoch (Tran et al. 1997, PASP 109, 489).

2003 June 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green